OpenStudy (anonymous):

What is the area of the largest rectangle with lower base on the x-axis and upper vertices on the curve y = 3 − x2

OpenStudy (anonymous):

did you draw a nice picture?

OpenStudy (anonymous):

i mean of \(y=3-x^2\) ?

OpenStudy (anonymous):

above the x axis only you know what it looks like?

OpenStudy (anonymous):

yup

OpenStudy (michele_laino):

I think that we have to maximize this function: \[\Large f\left( x \right) = 2x\left( {3 - {x^2}} \right)\] since we have this drawing: